Sample 1 (Contains His/leu)

RNA concentration 943.69 ng/µl. Purity 2.15 (OD 260/280)

Sample 2 (Contains 1a2a)

RNA concentration 217.96  ng/µl. Purity 2.13 (OD 260/280)

Sample 3 (Contains His/leu)

RNA concentration 1435.43 ng/µl. Purity 2.16 (OD 260/280)

Sample 4 (Contains 1a2a)

RNA concentration 1263.04 ng/µl. Purity 2.12 (OD 260/280)

Step 1. Math! You want to calculate how many µl will give 10 µg or 15 µg RNA. Your pre-lab question is (1) to compute the volume of RNA required for 10 or 15 µg for ALL 4 of the samples that your group* isolated during Lab 7. Then (2), determine the volume of water needed such that the volume of water and 10 (or 15) µg RNA is 44 µl. To clarify: if the sample contains His/Leu, calculate for 15 µg but if the sample contains 1a/2a, calculate for 10µg.
Answer #1

Sample 1, Concentration of RNA = 943.69 ng/l or 0.943g/l

Volume required for 15g RNA = 15/0.943 = 15.90l

Volume of water required = 44-15.90 = 28.1l

Sample 2, Concentration of RNA = 217.96ng/l or 0.217g/l

Volume required for 10g RNA = 10/0.217 = 46.08

Since the volume is exceeding 44l therfore water will not be added to this.

Sample 3, Conc of RNA = 1435.43ng/l or 1.43g/l

Volume required for 15g RNA = 15/1.43 = 10.48l

Volume of water required = 44-10.48 = 33.52l

Sample 4, Conc of RNA = 1263.04ng/l or 1.263g/l

Volume required for 10g RNA = 10/1.26 = 7.93l

Volume of water required = 44-7.93 = 36.07l